# Stable-V2A: Synthesis of Synchronized Sound Effects with Temporal and Semantic Controls

### v2

@@ Abstract
Sound designers and Foley artists usually sonorize a scene, such as from a movie or video game, by manually annotating and sonorizing each action of interest in the video. In our case, the intent is to leave full creative control to sound designers with a tool that allows them to bypass the more repetitive parts of their work, thus being able to focus on the creative aspects of sound production. We achieve this presenting Stable-V2A, a two-stage model consisting of: an RMS-Mapper that estimates an envelope representative of the audio characteristics associated with the input video; and Stable-Foley, a diffusion model based on Stable Audio Open that generates audio semantically and temporally aligned with the target video. Temporal alignment is guaranteed by the use of the envelope as a ControlNet input, while semantic alignment is achieved through the use of sound representations chosen by the designer as cross-attention conditioning of the diffusion process. We train and test our model on Greatest Hits, a dataset commonly used to evaluate V2A models. In addition, to test our model on a case study of interest, we introduce Walking The Maps, a dataset of videos extracted from video games depicting animated characters walking in different locations. Samples and code available on our demo page at https://ispamm.github.io/Stable-V2A.
@@
video-to-audio, sound effects synthesis, diffusion models, audio-video synchronization, multimodal audio synthesis.
@@ I Introduction
Sound plays a key role in audiovisual storytelling. A scene, whether from a movie, video game, or any other kind of audiovisual media, can acquire completely different meanings depending on how it is sonorized. To do so, Foley artists and sound designers meticulously research sounds that can provide the right emotional impact to the audience; as an example we can consider the difference that may exist in the footstep sounds in a horror movie, where a certain degree of suspense is required, and a comedy. This means that creatives must choose the desired sound to be used for the scene and then manually annotate all occurrences of the action to be sonorized. Such mechanical action can be time-consuming, not allowing the artist to focus on the most important part of audiovisual narration: the choice and processing of sounds.
@@
Figure 2: Stable-V2A architecture consists of two distinct parts: RMS-Mapper, that predicts an envelope representative for the audio directly from the input video, and Stable-Foley, the audio synthesis model for the controlled generation of the final audio effect. The generation is controlled temporally by the predicted RMS envelope through a DiT ControlNet, and semantically by CLAP and CAVP embeddings. The length of the output waveform can be controlled with $\mathrm{seconds\_start}$ and $\mathrm{seconds\_total}$ parameters.
Thanks to the latest advancements in deep learning and the flourishing of multimodal generative models [1, 2, 3] and LLMs [4, 5], research in the audio field is getting increasing attention, as it is now feasible to control sound generation through different modalities not strictly related to audio signal processing, such as text [6, 7, 8, 9, 10]. Many recent research papers have focused on developing models for the generation of sounds that can be used as a soundtrack for video or images, opening up a new branch of research in the audio field, commonly referred to as Video-to-Audio [11, 12, 13, 14, 15].
The definition of a Video-to-Audio (V2A) task refers to a whole class of problems dealing with the generation of audio that is semantically and temporally aligned with a reference input video. Early research work in this area focused on developing deep learning models that could generate an audio that is only semantically aligned with an input image [13, 16, 17] or video [12, 18]. The intent in these cases is to generate a soundtrack for the scene that would highlight its overall mood [19, 20]. Models of this kind can successfully generate musical accompaniment and ambient sounds [21, 22] that can be used as background track. However, when specific actions, such as footsteps or the clapping of hands, are to be sonorized, a precise temporal alignment is required in order to give the scene all the realism it needs: it only takes a few milliseconds of asynchronism between the reference video and the audio for the audience to perceive a feeling of estrangement [23].
Many papers that have attempted to achieve temporal as well as semantic alignment have focused on the design of end-to-end models that directly take as input the video to be sonorized and output the audio track related to that video [12, 15, 24, 25, 26, 27]. Those models have shown remarkable results in V2A tasks but do not allow sound designers or Foley artists to have direct control over the audio generation process, other than being able to directly operate on the final output of the model. Other works, however, have approached the problem by dividing it into two distinct components: a model for extracting audio-relevant features from the video and an audio model that using those features, generates the audio of interest [28, 29]. Our model is a two-stage model as well. In fact, as in our previous work [30], we aim to use a video model to extract human-readable features from the video that are representative of the audio to be generated; in this paper, such features are represented by the envelope, as the envelope has proven to be an excellent control for the temporal alignment [31].
We do so because our intent is to provide creatives with the ability to control the generation process. For example, the sound designer might decide to make some sounds more intense than normal or add other sounds that are not directly visible, such as the case where a person’s footsteps precede their visual presence on the scene. In this way, by extracting a feature like the envelope from a video, the sound designer can directly edit this envelope to his convenience, before using it as a control for generating the final audio track.
Here we introduce a new two-stage model for V2A generation, Stable-V2A, where the two stages are represented by RMS-Mapper and Stable-Foley. The architecture of our complete model is shown in Fig. 2
RMS-Mapper is a simple but effective network for mapping an RMS envelope directly from a video; it takes representative features of frames and optical flow of videos as input. We use a widely employed video encoder for audiovisual tasks [32] as baseline. The output is a curve that aims to replicate the RMS envelope, in terms of time alignment, time-varying intensity, and duration of each sound present in the target audio.
Stable-Foley, the model for synthesizing sound effects, relies on Stable Audio [33], a state-of-the-art latent diffusion model for audio generation, to obtain the final output. In order to take advantage of the prior knowledge of this model and guide its generation process via the RMS envelope, we leverage a ControlNet [34], which is a network used to control the diffusion process via extra conditioning.
Fig. 1 depicts an audio generated by Stable-Foley through the RMS-envelope predicted from RMS-Mapper.
The semantics of the final audio is specified by conditioning Stable-Foley with a latent representation of the audio sample representing the sound of the action in the video. This representation is obtained through CLAP [35], a state-of-the-art audio encoder based on contrastive learning, that aligns audio and textual modalities. Therefore, the desired timbre for the final audio track can be specified to the model either with a reference audio sample as well as with a textual description of the sound, also allowing for text-to-audio generation. Our model generates 44.1 kHz stereo audio of variable length.
We evaluate the proposed method on several objective metrics by training and testing the model on Greatest Hits [36], a dataset widely used to validate V2A models, consisting of videos of people hitting or scratching objects with a drumstick.
Finally, to test the model on a case study of interest, which is the generation of footstep sounds [37], we introduce a dataset of videos extracted from video games session containing animated characters walking in different environments, called Walking The Maps. The videos are taken from YouTube and selected in order to have as few background noises as possible, so that the audio of each video contains mainly the sound of footsteps.
Thus, our contributions can be summarized as follows:
  • We present Stable-V2A, a two-stage model for generating audio that is temporally and semantically aligned with a reference input video. Sound designers can control the generation of the final output through the use of an RMS envelope.
  • We perform an evaluation through objective metrics that attest both temporal and semantic alignment capability of the model.
  • We also introduce a brand-new dataset for the analysis of a case study of interest, which is the generation of footstep sounds, called Walking The Maps.
@@ II Related Works
Generating audio that is aligned with a silent reference video is a problem of great interest today for multimedia post production. Due to recent developments in deep learning, many of the leading experts in the audiovisual industry are focusing their efforts on developing deep learning models that can be integrated into the principal post production tools [38, 39, 40, 41].
@@
Im2Wav [13] focuses on generating sounds that are semantically relevant to an image or sequence of images, using CLIP [46] to condition with visual features two transformer language models.
@@
In order to perform a cross-modality generation between video and audio, Xing, et al. [47] propose a multimodality latent aligner with the pre-trained ImageBind [48] model, used to condition a latent diffusion model (LDM) in order to generate an audio that is semantically relevant for an input video.
@@
RegNet [32] represents one of the first attempts to achieve semantic as well as temporal alignment. This model uses a simple and efficient video encoder that extracts relevant visual features from the frames and optical flow of an input video and uses them to condition a Generative Adversarial Network (GAN) to generate visually aligned sounds. SpecVQGAN [49] also uses RGB and optical flow features of a video but leveraging a more powerful Transformer-based autoregressive architecture to generate temporally and semantically aligned sounds to an input video.
@@
A similar architecture is used in CondFoleyGen [50], where such a model is trained directly using Greatest Hits, succeeding in achieving an efficient alignment in both content and timing with the reference video. Instead, Diff-Foley [51] uses Contrastive Audio-Visual Pretraining (CAVP) to temporally and semantically align audio and video modalities, being able to generate a video embedding that contains features relevant to the corresponding audio. Such features are used to directly condition an LDM.
However, all these models do not provide human-intelligible control, thus not allowing direct control by sound designers over the final generation, as they cannot act on either the semantics or the timing of the final output.
@@
Our previous work, SyncFusion [30] - which was the first to introduce a human readable control for the V2A task - uses a video encoder based on a ResNet(2+1)D-18 [52] that, by taking frames of the target video as input, generates an onset track; this onset track is then passed to a time-domain diffusion model [53] to generate the final output. The visual representation provided by the onsets is very similar to the manual annotations used by sound designers when determining the temporal placements of the sound sources to be sonorized in a video. Therefore, it provides an informative and yet easy-to-edit control for the user, allowing direct oversight on the generation of the final output. This onset track is represented by a binary mask that indicates the presence or absence of the action of interest for each frame of the video. Hence, an onset can indicate the temporal location of a sound but can not inform either about the intensity or temporal duration.
Furthermore, the onset of a sound event cannot be determined unambiguously, as in the case of a sound event sustained over time where the temporal placement of the onset is not unequivocally determined. Consequently, in order to use onsets, it is necessary to have a dataset containing manually annotated onsets for every sound event in each video.
@@ III-A Audio Diffusion Models
Denoising Diffusion Probabilistic Models were introduced in [55]. They were first used to obtain state-of-the-art results in image generation and then in the generation of other kind of data, such as video and audio.
Like Diffusion Models for images, Audio Diffusion Models (ADM) start by producing a random, noisy audio signal and gradually enhance it through multiple refining iterative steps. In each step, the noise is reduced, and finer details are added to the audio. Such models have been used to generate both music [56] and environmental sounds [57].
@@
These models are widely used for audio generation [59], with Stable Audio [60, 61] representing a state-of-the-art latent ADM in generative deep learning for audio.
@@ III-B Guiding diffusion processes with ControlNet
ControlNet [34] is a neural network specifically designed for diffusion models that allows for more precise control over the generation process by conditioning the model on additional inputs. It was first used to control image generation in Stable Diffusion with extra information such as features or prompts. Initially designed for U-Net-based diffusion models, ControlNet was later extended in [62] to work with Diffusion Transformers (DiT).
@@ IV-A Problem Formulation
Let us consider a target video $\mathbf{x}\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ is the total duration of the video expressed in frames, $C$ is the number of input RGB channels, i.e. 3 channels, $H\times W$ is the dimension in height and width of each frame. The objective of the proposed model is to generate an audio track $\mathbf{y}\in\mathbb{R}^{Ch\times L}$, where $Ch$ represents the number of audio channels, that is 2 for stereo audio, and L is the time duration of the audio expressed in samples. The generated audio $\mathbf{y}$ must be semantically and temporally aligned to the input video, so it can be used as a realistic soundtrack for it. For example, in the case of a video depicting a person walking slowly on a wooden pavement, the audio generated by the model must be both in terms of semantics and temporal alignment distant from the audio generated for a video depicting a person running on grass.
One of the main challenges when approaching V2A problems is the completely different temporal resolution between video and audio. Indeed, the temporal resolution of a video, expressed in frames per second (fps), is usually much lower than the temporal resolution of an audio, expressed in sample rate (sr). For instance, a commonly used frame rate in video is 25 fps, while the standard in audio is represented by a sr of 44.1 kHz. This means that in this specific case 1764 audio samples are represented by one single frame. One possible solution is to increase the fps of the video and lower the sr of the audio, to mitigate that difference, as we partially do in our video model. However, increasing the frame rate of a video excessively leads to a significant rise in the computational cost of the model. On the other hand,, decreasing the sample rate results in a loss of audio quality, which is unsuitable for practical use in sound design. It is therefore crucial to carefully manage the transition between the two modalities, i.e. audio and video, to ensure the problem remains computationally feasible while maintaining the audio quality necessary to capture all the sound features needed by sound designers and Foley artists [64, 65].
@@ IV-B Mapping a Video to an Envelope
Following what was first done in [54], we address this task as a classification one since trying to map a curve representing the envelope of an audio as a regression task does not produce satisfactory results. This is because sounds related to actions are often transient, which makes audio containing multiple repeated actions sparse, meaning that most samples represent silence. A model that tries to map such a time series as a regression task thus tends to predict silence or the mean value of the curve, as we also found out in our early experiments.
We then perform amplitude quantization of the envelope by $\mu$-law encoding. Such encoding is widely used in audio, especially in pitch estimation [66], precisely to mitigate the transient and sparse nature of waveform of this kind. The $\mu$-law encoding reduces the dynamic range of an audio signal and is formulated as follows:
$\mu-\mathrm{law}(\mathbf{r})=\frac{\ln(1+\mu|\mathbf{r}|)}{\ln(1+\mu)}\cdot\text{sgn}(\mathbf{r}),$ (8)
where $\mathbf{r}$ is the RMS envelope of the target audio $\mathbf{y}$.
@@
This encoding allows amplitude quantization of the signal by dividing the values into equidistant bins. The idea here is to map video frames to specific classes, representing the amplitude values of the envelope. This conceptually is very close to what we did in SyncFusion [30], by simply extending the binary classification task of an onset track to the multi-class classification task, where each class represents a given range of the continuous envelope amplitude values.
Furthermore, it is necessary not to penalize near-correct predictions. Indeed, a precise prediction of the correct class is desirable but not strictly required to ensure good perceptual quality: by using 64 equidistant bins as classes, each would correspond to an amplitude resolution of 0.50 dB. Selecting a bin next to the true one would introduce an error that is barely perceptible.
Then, in order to minimize the penalty of a near-correct class prediction, Gaussian label smoothing on the target envelopes - as proposed in [67] - is used:
$\mathbf{r}(i)=\begin{cases}\exp\left(-\frac{(c_{i}-c_{\text{gt}})^{2}}{2\sigma^{2}}\right)&\text{if }|c_{i}-c_{\text{gt}}|\leq W\,(c_{i},c_{\text{gt}}\neq 0)\\ 0&\text{otherwise},\end{cases}$

### v3

# with Semantic and Temporal Alignment

@@ Abstract
Traditional sound design workflows rely on manual alignment of audio events to visual cues, as in Foley sound design, where everyday actions like footsteps or object interactions are recreated to match the on-screen motion. This process is time-consuming, difficult to scale, and lacks automation tools that preserve creative intent. Despite recent advances in vision-to-audio generation, producing temporally coherent and semantically controllable sound effects from video remains a major challenge. To address these limitations, we introduce Fol·AI, a two-stage generative framework that decouples the when and the what of sound synthesis, i.e., the temporal structure extraction and the semantically guided generation, respectively. In the first stage, we estimate a smooth control signal from the video that captures the motion intensity and rhythmic structure over time, serving as a temporal scaffold for the audio. In the second stage, a diffusion-based generative model produces sound effects conditioned both on this temporal envelope and on high-level semantic embeddings, provided by the user, that define the desired auditory content (e.g., material or action type). This modular design enables precise control over both timing and timbre, streamlining repetitive tasks while preserving creative flexibility in professional Foley workflows. Results on diverse visual contexts, such as footstep generation and action-specific sonorization, demonstrate that our model reliably produces audio that is temporally aligned with visual motion, semantically consistent with user intent, and perceptually realistic. These findings highlight the potential of Fol·AI as a controllable and modular solution for scalable, high-quality Foley sound synthesis in professional and interactive settings. Supplementary materials, including samples and code, are accessible on our dedicated demo page at https://ispamm.github.io/FolAI.
@@
video-to-audio, sound effects synthesis, diffusion models, audio-video synchronization.
@@ I Introduction
Sound is a fundamental component of audiovisual storytelling, capable of deeply altering the meaning and emotional impact of a scene. Traditionally, Foley artists and sound designers manually select, record, and synchronize sounds (like footsteps or object interactions) to match on-screen actions. This careful process ensures the desired narrative effect but is inherently time-consuming and requires significant effort. Recent advancements in deep learning, particularly in multimodal generative models [1, 2, 3] and Large Language Models (LLMs) [4, 5], have generated significant interest in automating and enhancing audio synthesis. This has led to the rise of Video-to-Audio (V2A) generation [6, 7, 8, 9, 10], a research area focused on synthesizing audio that is semantically relevant and temporally aligned with an input video sequence. Initial efforts in this area often focused on semantic alignment, generating ambient sounds or soundtracks reflecting the overall mood [11, 12, 13, 14], but lacked the temporal precision needed for synchronizing specific sound effects.
Achieving both appropriate semantic content and precise timing is crucial for effective V2A generation as temporal alignment is equally fundamental for realism, particularly when synchronizing distinct sound events like Foley effects. Even slight timing mismatches, on the order of milliseconds, can be perceived by the audience and disrupt the feeling of immersion [15]. Furthermore, although more recent end-to-end V2A models [7, 10, 16, 17, 18, 19] aim for better synchrony, they often function as “black boxes”. This typically offers limited direct control for sound designers to refine the timing or creatively adjust the generated audio beyond basic post-processing. This lack of fine-grained control hinders their adoption in professional workflows where artistic input and adjustments are essential.
To address these critical needs for robust temporal alignment and user control, we introduce FOL·AI, a novel two-stage generative framework designed for accurate V2A synthesis. Our approach separates the task into predicting the temporal structure and then generating the audio based on that structure and desired sound characteristics, as shown in Fig. 2. The first stage analyzes the input video using frame features and optical flow to predict a Root Mean Square (RMS) envelope. This envelope represents the audio’s timing, intensity, and duration [20]. Importantly, this predicted envelope is human-readable and can be directly edited by the sound designer. This allows for fine-tuning the timing or even adding sounds for off-screen events before the audio is generated. The second stage employs a state-of-the-art diffusion model architecture to create the final sound. To ensure the audio matches the timing specified by the envelope, we introduce an additional conditioning network architecture that steers the diffusion process using the RMS envelope from the first stage. Users can control the timbre by providing semantic information through embeddings from reference audio samples, text descriptions and the input video. This modular design allows creatives to control when sounds occur and what they sound like.
The FOL·AI framework successfully generates 44.1 kHz stereo audio of variable length that shows strong temporal and semantic synchronization with the input video, as qualitatively illustrated in Fig. 1. We demonstrate the effectiveness of our model through careful objective evaluations on the widely used Greatest Hits V2A benchmark dataset [21], where FOL·AI achieves state-of-the-art results compared to existing V2A models. Furthermore, recognizing the lack of suitable public datasets with high-quality audio and video for evaluating V2A models on specific, highly relevant sound design tasks like Foley, we introduce Walking The Maps. This new dataset focuses specifically on footstep generation, a key challenge and frequent task in the field [22]. Sourced from high-quality video game walkthroughs available online, it provides numerous video clips featuring clearly audible footstep sounds across various surfaces (like grass, concrete, wood) and character actions (walking, running), paired with corresponding high-definition video offering excellent temporal cues. The audio was carefully preprocessed using source separation techniques to isolate footstep sounds, creating clean targets suitable for robust V2A model training and evaluation. We make Walking The Maps freely available to the research community to facilitate progress and benchmarking in realistic automated Foley synthesis scenarios.
@@
Figure 2: FOL·AI architecture consists of two distinct parts: the video model, that predicts an envelope representative for the audio directly from the input video, and the audio synthesis model for the controlled generation of the final audio effect. The generation is controlled temporally by the predicted RMS envelope through a DiT ControlNet, and semantically by CLAP and CAVP embeddings. The length of the output waveform can be controlled with $\mathrm{seconds\_start}$ and $\mathrm{seconds\_total}$ parameters.
@@ II Related Works
Generating audio that is aligned with a silent reference video is salient task nowadays for multimedia post production. Due to recent developments in deep learning, many of the leading experts in the audiovisual industry are focusing their efforts on adopting deep learning models that can be integrated into the principal post production tools [23, 24, 25, 26].
@@
Im2Wav [8] focuses on generating sounds that are semantically relevant to an image or sequence of images, using CLIP [31] to condition with visual features two transformer language models. In order to perform a cross-modality generation between video and audio, Xing, et al. [32] propose a multimodality latent aligner with the pre-trained ImageBind [33] model, used to condition a latent diffusion model (LDM) in order to generate an audio that is semantically relevant for an input video.
@@
RegNet [34] represents one of the first attempts to achieve semantic as well as temporal synchronization. This model uses a simple and efficient video encoder that extracts relevant visual features from frames and optical flow of an input video and uses them to condition a Generative Adversarial Network (GAN) to generate visually aligned sounds. SpecVQGAN [35] also uses RGB and optical flow features of a video but leveraging a more powerful Transformer-based autoregressive architecture to generate temporally and semantically synchronized sounds to an input video.
@@
A similar architecture is used in CondFoleyGen [36], where such a model is trained directly using Greatest Hits, succeeding in achieving an efficient alignment in both content and timing with the reference video. Instead, Diff-Foley [37] uses Contrastive Audio-Visual Pretraining (CAVP) to temporally and semantically align audio and video modalities, being able to generate a video embedding that contains features relevant to the corresponding audio. Such features are used to directly condition an LDM. However, all these models do not provide human-intelligible control, thus not allowing direct supervision by sound designers over the final generation, as they cannot act on either the semantics or the timing of the final output.
@@
SyncFusion [38] - which was the first to introduce a human readable control for the V2A task - uses a video encoder based on a ResNet(2+1)D-18 [39] that, by taking frames of the target video as input, generates an onset track; this onset track is then passed to a time-domain diffusion model [40] to generate the final output. A similar approach is implemented in [41]. The visual representation provided by the onsets is very similar to the manual annotations used by sound designers when determining the temporal placements of the sound sources to be sonorized in a video. Therefore, it provides an informative and yet easy-to-edit control for the user, allowing direct oversight on the generation of the final output. This onset track is represented by a binary mask that indicates the presence or absence of the action of interest for each frame of the video. Hence, an onset can indicate the temporal location of a sound but can not inform either about the intensity or temporal duration. Furthermore, the onset of a sound event cannot be determined unambiguously, as in the case of a sound event sustained over time where the temporal placement of the onset is not unequivocally determined. Consequently, in order to use onsets, it is necessary to have a dataset containing manual annotation for every sound event in each video.
@@ III-A Audio Diffusion Models
Denoising Diffusion Probabilistic Models were introduced in [44]. They were first used to obtain state-of-the-art results in image generation and then in the synthesis of different media, such as video and audio. Like Diffusion Models for images, Audio Diffusion Models (ADM) start by producing a random, noisy audio signal and gradually enhance it through multiple refining iterative steps. In each step, the noise is reduced, and finer details are added to the audio. Such models have been used to generate both music [45] and environmental sounds [46].
@@
These models are widely used for audio generation [48], with Stable Audio [49, 50] representing a state-of-the-art latent ADM.
@@ III-B Guiding diffusion processes with ControlNet
ControlNet [51] is a neural network specifically designed for diffusion models that allows more precise control over the generation process by conditioning the model on additional inputs. It was first used to control image generation in Stable Diffusion with extra information such as features or prompts. Initially designed for U-Net-based diffusion models, ControlNet was later extended in [52] to work with Diffusion Transformers (DiT) for image synthesis.
@@ IV-A Problem Formulation
Let us consider a target video $\mathbf{x}\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ is the total duration of the video expressed in frames, $C$ is the number of input RGB channels, i.e. 3 channels, $H\times W$ is the dimension in height and width of each frame. The objective of the proposed model is to generate an audio track $\mathbf{y}\in\mathbb{R}^{Ch\times L}$, where $Ch$ represents the number of audio channels, that is 2 for stereo audio, and L is the time duration of the audio expressed in samples. The generated audio $\mathbf{y}$ must be semantically and temporally synchronized with the input video, so it can be used as a realistic soundtrack for it. For example, in the case of a video depicting a person walking slowly on a wooden pavement, the audio generated by the model must be both in terms of semantics and temporal alignment distant from the audio generated for a video depicting a person running on grass.
One of the main challenges when approaching V2A problems is the completely different temporal resolution between video and audio. Indeed, the temporal resolution of a video, expressed in frames per second (fps), is usually much lower than the temporal resolution of an audio, expressed in sample rate (sr). For instance, a commonly used frame rate in video is 30 fps, while the standard in audio is represented by a sr of 44.1 kHz. This means that in this specific case 1470 audio samples are represented by one single frame. One possible solution to mitigate this difference is to increase the fps of the video and lower the sr of the audio. However, increasing the frame rate of a video excessively leads to a significant rise in the computational cost of the model. On the other hand, decreasing the sample rate results in a loss of audio quality, which is unsuitable for practical use in sound design. It is therefore crucial to carefully manage the transition between the two modalities, i.e., audio and video, to ensure that the problem remains computationally feasible while maintaining high enough audio quality to capture all the features of the sound needed by sound designers and Foley artists [54, 55].
@@ IV-B Mapping a Video to an Envelope
We address this task as a classification one since trying to map a curve representing the envelope of an audio as a regression task does not produce satisfactory results [42]. The reason is that sounds related to actions are often transient, which makes audio containing multiple repeated actions sparse, i.e. most samples represent silence. A model that tries to map such a time series as a regression task thus tends to predict silence or the mean value of the curve, as we also found out in our early experiments.
We then perform amplitude quantization of the envelope by $\mu$-law encoding. This encoding is widely used in audio, especially in pitch estimation [56], precisely to mitigate the transient and sparse nature of waveform of this kind. This encoding allows amplitude quantization of the signal by dividing the values into equidistant bins. The idea here is to map video frames to a sequence of specific classes, representing the amplitude values of the envelope.
@@
Furthermore, it is necessary not to penalize near-correct predictions. Indeed, a precise prediction of the correct class is desirable but not strictly required to ensure good perceptual quality. Then, in order to minimize the penalty of a near-correct class prediction, Gaussian label smoothing on the target envelopes - as proposed in [57] - is used:
$\mathbf{r}(i)=\begin{cases}\exp\left(-\frac{(c_{i}-c_{\text{gt}})^{2}}{2\sigma^{2}}\right)&\text{if }|c_{i}-c_{\text{gt}}|\leq W\,(c_{i},c_{\text{gt}}\neq 0)\\ 0&\text{otherwise},\end{cases}$ (8)